find the points of intersection, if any, of the graphs in the system.
x^2+y^2=17
y=x+3

Answer by lwsshak3(11628)

find the points of intersection, if any, of the graphs in the system.
x^2+y^2=17
y=x+3
***
sub(x+3) for y in first equation
x^2+(x+3^2=17
x^2+x^2+6x+9=17
2x^2+6x-8=0
x^2+3x-4=0
(x+4)(x-1)=0
..
x=-4
y=x+3=-1
point of intersection: (-4,-1)
..
x=1
y=x+3=4
2nd point of intersection: (1,4)